LASER SPECTROSCOPY OF THE B^{2}Σ^{+} − X^{2}Σ^{+} TRANSITION OF ZrN AND TiN HONGBING CHEN, YONGFANG LI, and \underline{A. S-C. CHEUNG}, Department of Chemistry, The University of Hong Kong, Pokfulam Road, Hong Kong.